A man is q years old while his son is p years old.the sum of their age is equal to twice the difference of their ages.the product of the age is 675.write down the equation connecting their age and solve the equation in order to find the ages of the man and his son?
Answer by Dickiedannie(6)

ANSWER::::
i. P+Q=2(P-Q), P*Q=675
P+Q=2P-2Q
"3Q=P" EQUATION CONNECTING THEIR AGES
ii.BUT WE ARE GIVEN P*Q=675, THEN INPUT P=3Q INTO P*Q=675 =>3Q^2=675 SOLVE FOR Q AND GET Q=15, FROM 3Q=P THEN 3*15=P=45

THEREFORE, THE EQUATION CONNECTING THEIR AGES (MAN AND HIS SON) IS (P=3Q) AND THEIR AGES ARE 45 AND 15 YEARS RESPECTIVELY.